OpenStudy (anonymous):

Find the inverse of the function. f(x) = 6x^3 - 3

OpenStudy (anonymous):

First, switch the X and Y, x=6y^3-3 Then, solve for y x+3=6y^3 (x+3)^(1/3)=6y ((x+3)^(1/3))/6=y

OpenStudy (anonymous):

B